Four identical 5.00 Ω resistors are joined and form the four sides of a square. (a) What is the resistance between points A and B? (b) What is the resistance between A and C?
Answer #1

a) between A and B, three resistors are in series and they are paralle to the fourth resistor

Resistnace = R

1/R=1/5+1/15

R=3.75 Ohms

b)between A and C, two 5 Ohm resistors iare are series in each brace and each brabch is parallel

1/R=1/10 + 1/10

R=5 Ohms

Resistnace between A and C is 5 Ohms